The longer leg of a right triangle is 1 inch longer than the shorter leg. The hypotenuse is 9 inches longer than the shorter leg. Find the side length of the triangle.
Answer by jorel555(1290)

Let n be the length of the shortest leg. Then the other leg is n+1; and the hypotenuse is n+9. So:
n²+(n+1)²=(n+9)²
2n²+2n+1=n²+18n+81
n²-16n-80=0
(n-20)(n+4)=0
n=20 or -4
Throwing out the negative result, we get the length of the shorter leg to be 20 inches, the longer leg to be 21 inches, and the hypotenuse is 29 inches. ☺☺☺☺